If a car covers 2/5th otlf the total distance with vl speed and the 3/5th distance with v2 then average speed is ____
1
Expert's answer
2021-03-08T08:17:32-0500
"v_{av}=\frac{s}{t}\\v_{av}=\frac{s}{\frac{2s}{5v_1}+\frac{3s}{5v_2}}\\
\\v_{av}=\frac{5v_1v_2}{3v_1+2v_2}"